OpenStudy (anonymous):

On which of the following lines does the point (1, 6) lie? a. y = x + 5 c. y = 2x - 1 b. y = -x + 7 d. both a and b

OpenStudy (zale101):

plug in the numbers to each equation (from the choices) and see if it matches the point

OpenStudy (zale101):

a. y=x+5 6=1+5 6=6 correct c. y=2x-1 6=2(1)-1 6=2-1 6=1? no b. y = -x + 7 6=-(1)+7 6=6 Correct!